OpenStudy (anonymous):

Total number of DVDs produced in 1998 was0.5 million. in 2004 the total number was 29.01 million. assuming an exponential model applies a) find the value of k and write the function b) estimate # of dvds in 2005,2008, and 2011

OpenStudy (anonymous):

x(t) = x0 exp(k*t) x0 = x(0) = 0.5*10^6 x(6) = x(0) exp(k*6) x(6)/x(0) = exp(k*6) k = 1/6*log(x(6)/x(0)) = 1/6*log(29.01/0.5) For 2005, t = 7 x(7) = x0 exp(k*7) for 2008, t = 10 x(10) = x0*exp(k*10) for 2011, t = 13 x(13) = x0*exp(k*13)